OpenStudy (anonymous):

Fan and Medal for best answer! Ease show how you got it. Solve by Completing the Square. x^2+x=30

OpenStudy (anonymous):

half of 1 is \(\frac{1}{2}\) and \((\frac{1}{2})^2=\frac{1}{4}\) so start with \[(x+\frac{1}{2})^2=30+\frac{1}{4}\]

OpenStudy (anonymous):

did that